### v2

# Are ELECTRA’s Sentence Embeddings Beyond Repair? The Case of Semantic Textual Similarity

Abstract
While BERT produces high-quality sentence embeddings, its pre-training computational cost is a significant drawback. In contrast, ELECTRA delivers a cost-effective pre-training objective and downstream task performance improvements, but not as performant sentence embeddings. The community tacitly stopped utilizing ELECTRA’s sentence embeddings for semantic textual similarity (STS). We notice a significant drop in performance when using the ELECTRA discriminator’s last layer in comparison to earlier layers. We explore this drop and devise a way to repair ELECTRA’s embeddings, proposing a novel truncated model fine-tuning (TMFT) method. TMFT improves the Spearman correlation coefficient by over points while increasing parameter efficiency on the STS Benchmark. We extend our analysis to various model sizes, languages, and two other tasks. Further, we discover the surprising efficacy of ELECTRA’s generator model, which performs on par with BERT, using significantly fewer parameters and a substantially smaller embedding size. Finally, we observe further boosts by combining TMFT with a word similarity task or domain adaptive pre-training.
Are ELECTRA’s Sentence Embeddings Beyond Repair?
The Case of Semantic Textual Similarity
Ivan Rep David Dukić Jan Šnajder TakeLab, Faculty of Electrical Engineering and Computing, University of Zagreb irep2718@gmail.com {david.dukic, jan.snajder}@fer.hr
1 Introduction
Pre-trained language models (PLMs) have been a staple in NLP for years, leveraging self-supervised objectives to improve representations for downstream tasks. BERT (Devlin et al., 2019), one of the most widely used PLMs, uses a masked language modeling (MLM) objective for pre-training. The main drawbacks of MLM are the substantial compute cost due to the low masking rate, and the gap between the pre-training task and downstream tasks. To address these defects, Clark et al. (2020) introduce ELECTRA, which substitutes MLM with replaced token detection (RTD), achieving the same results as BERT but with four times less compute. RTD uses a generator model that corrupts the input, while a discriminator model distinguishes between corrupted and original tokens.
Semantic textual similarity (STS) Agirre et al. (2013) is a foundational NLP task with broad applications. STS applications must balance accuracy and embedding size to ensure fast inference. To this end, Reimers and Gurevych (2019) introduced sentence transformers, a framework based on the transformer bi-encoder architecture. It encodes text representations independently before pooling them, calculating the similarity score through a comparison operation between the two embedded texts.
While some transformer models excel across GLUE (Wang et al., 2018) tasks, their sentence embedding quality may vary. Sentence transformers often use PLMs pre-trained with a language modeling objective (Song et al., 2020; Raffel et al., 2020; Liu et al., 2019). In contrast, RTD pre-trained models demonstrate subpar performance in a bi-encoder setting (Reimers, 2021), but pre-training such a model is more cost-effective. It is unclear whether this low performance is due to pre-training and whether the embeddings can be improved.
To address this gap, we present a layer-wise study of ELECTRA’s sentence embeddings. We examine the performance discrepancy between BERT and ELECTRA’s discriminator across various model sizes, languages, and tasks. We hypothesize the last few layers of the discriminator are too specialized for the pre-training task. Following this, we devise truncated model fine-tuning (TMFT), shown in Figure 1, which utilizes the transformer up to layer , followed by pooling and fine-tuning the whole truncated model. We show that the discriminator suffers from a drop in STS performance when fine-tuning the final layer’s embeddings and that this drop is consistent across all inspected model sizes and languages. Although we focus on STS, we expose a similar trend for two other tasks. When applying TMFT on STS Benchmark (STSB) (Cer et al., 2017), the TMFT-obtained embeddings significantly outperform the ones from the ELECTRA discriminator’s last layer. Moreover, TMFT-obtained embeddings outperform the ones of BERT fine-tuned on STSB up to but not including the eighth layer. Further, we uncover the efficacy of the ELECTRA generator model, performing on par with BERT while having a significantly smaller embedding size and substantially fewer parameters. Finally, we propose two improvements of the basic TMFT method: prior TMFT on word similarity (WS) and prior domain adaptive pre-training (DAPT) using MLM (cf. Figure 1).
Our contributions are: (1) a layer-wise analysis of ELECTRA’s sentence embeddings for various model sizes, languages, and tasks; (2) a novel TMFT method for repairing ELECTRA’s embeddings that substantially improves performance on STS, paraphrase identification (PI), and textual entailment (TE) tasks, and uncovering the surprising efficacy of the generator model; (3) two additional techniques in combination with TMFT for improving ELECTRA’s embeddings even further on STS.
2 Related Work
Reimers and Gurevych (2019) introduced Siamese networks to transformers, sparking significant research aimed at enhancing sentence embeddings. Fine-tuning on an auxiliary task unrelated to STS has also been explored. Reimers and Gurevych (2019) first fine-tune on a TE task, followed by fine-tuning on STS. Furthermore, DAPT has become a widely adopted method, enhancing performance in downstream tasks (Gururangan et al., 2020).
Similarly, there has been considerable interest in using different layers of a PLM for sentence embeddings. Bommasani et al. (2020) assess the layer-wise performance of various transformer models and pooling methods on WS datasets. Huang et al. (2021) determine what combinations of hidden states perform the best for unsupervised STS, while Jawahar et al. (2019) extract layer-wise structural characteristics encoded in BERT with probing tasks. Finally, Ethayarajh (2019) explores layer-wise embedding anisotropy. Two works that resemble ours the most are Hosseini et al. (2023) and Li et al. (2024). The former improves results by combining layer representations with dynamic programming, whereas we fine-tune the model from input embeddings up to a specific layer. The latter, developed concurrently with our method, uses model and embedding truncation combined with fine-tuning and a specialized loss function.
3 Truncated Model Fine-Tuning
The usual approach for obtaining sentence embeddings is applying a pooling operation over the last layer’s embeddings. We use mean pooling as it yields the best results, in line with previous work (Reimers and Gurevych, 2019). The TMFT method we propose for repairing ELECTRA’s embeddings reduces to taking the -th layer output followed by pooling and fine-tuning on the target task.
For a sentence the encoder outputs a tensor , where is the number of layers including the input embeddings (which we treat as a layer). We then apply mean pooling over the -th representation, , where is the -dimensional output token embedding of layer for token at position . We apply this to both sentences, compare them using cosine similarity, and propagate the loss from layer to the model input.111Fine-tuning based on the final representation and using prior representations for inference does not yield satisfying results, and a similar approach has already been tried out (Hosseini et al., 2023).
Furthermore, we propose combining one supervised and one self-supervised method with TMFT for performance gains: (1) PLM fine-tuning on the WS task or (2) DAPT using MLM. The intuition for the former is that WS is crucial for assessing sentence similarity, while the latter builds upon prior evidence that MLM-based PLMs work well in a bi-encoder setting (Reimers, 2021).
4 Experiments and Results
We fine-tune each model on the STSB from GLUE. We also run the experiments on machine-translated versions of STSB in Korean, German, and Spanish (cf. Appendix B). We choose these based on language-specific dataset and PLM availability. We report the Spearman’s correlation coefficient (SCC) as suggested by Reimers and Gurevych (2019). For WS experiments, we use the set of all pairs present in the following datasets: RG- (Rubenstein and Goodenough, 1965), WordSim- (Finkelstein et al., 2001), SimLex- (Hill et al., 2015), and SimVerb- (Gerz et al., 2016), with a random :: train, validation, and test split. The experiments with MLM are conducted on a dataset of sentences from SNLI (Bowman et al., 2015) and MultiNLI (Williams et al., 2018). We extend our experiments to PI and TE tasks using the Microsoft Research Paraphrase Corpus (MRPC) from GLUE and Sentences Involving Compositional Knowledge (SICK) entailment dataset (Marelli et al., 2014), respectively (cf. Appendix A). For STSB, MRPC, and SICK, we use the cross-validation splits defined by the dataset authors.
We use the following models from HuggingFace Transformers (Wolf et al., 2020): BERT tiny, mini, small, medium, base, and large. For the ELECTRA discriminator and generator, we use the small, base, and large models. To strengthen our findings on RTD-pre-trained models, we run experiments for the DeBERTaV3 model (He et al., 2021). For fine-tuning and pre-training, we use the AdamW optimizer Loshchilov and Hutter (2017), a learning rate of , and weight decay set to . We apply gradient clipping to a max norm of . For fine-tuning on all tasks except WS, we use a batch size of for epochs. WS fine-tuning uses a batch size of for epochs. For DAPT, we use a batch size of with gradient accumulation steps for epochs with masking probability. All reported results are averaged across five seeds. Test set SCCs on STSB correspond to the model with the highest SCC on the validation set. For MRPC and SICK, the F1 scores correspond to the models with the classification threshold optimized for the highest validation set F1 score.
4.1 ELECTRA
Applying ELECTRA to downstream tasks is usually done using the discriminator. We decided not to follow this practice, as ELECTRA’s authors do not give convincing reasons for discarding the generator. Hence, we conduct experiments with both models. The generator is similar to BERT, except the generator’s input embeddings are tied to the discriminator in pre-training. For comparison, we use BERT as a baseline as it is a standard choice and similar in size to the discriminator.
Figure 2(a) shows test set SCCs for TMFT applied to ELECTRA discriminator, ELECTRA generator, and BERT. BERT shows a trend where the SCC almost monotonously increases as the index of the fine-tuned layer embedding increases. The same trend is present for the generator. ELECTRA generator with M parameters maintains comparable performance to BERT with M parameters on all tasks (cf. Table 1, Figure 4, and Figure 5 in Appendix D). This finding is consistent across all inspected generator and BERT sizes (cf. Table 2 in Appendix D). The discriminator shows a different trend, gradually increasing until the ninth layer, after which performance drops sharply. We attribute this drop to the RTD pre-training task, as confirmed by Centered Kernel Alignment (CKA) representation similarity analysis, which shows that CKA values between discriminator models and MLMs drop in the final layers even before fine-tuning (cf. Appendix E). A similar performance drop occurs for Korean, German, and Spanish (cf. Appendix B), all considered discriminator model sizes (cf. Figure 10), the PI task (cf. Figure 4), and the DeBERTaV3 discriminator (cf. Figure 2(a)). The increase in the ELECTRA discriminator test SCCs between the ninth (M parameters) and last state (M parameters) is points. The best-performing state on the validation set is the third state (M parameters), and the increase compared to the last state on the test set is . Furthermore, the discriminator’s input embeddings outperform BERT, while its output embeddings surpass BERT’s up to, but not including, the eighth layer. Across all models, the largest increase between consecutive layers is between the 0th and 1st layers, likely due to self-attention.
4.2 Further Improvements
The first improvement we propose is TMFT on WS before TMFT on sentence similarity. The downside of this method is that it requires data labeled for WS. We consider only using the same layer for both fine-tuning procedures. Figure 2(b) gives the test set SCCs on STSB for the proposed improvement. We observe BERT and the discriminator both benefit from this method, with ELECTRA outperforming BERT up to but not including the tenth layer.
Our second improvement is DAPT using MLM before TMFT on STS, regardless of the model type. We only consider pre-training using the last layer’s embeddings. Figure 2(c) presents the test set SCCs on the STSB dataset. Compared to excluding improvements, the performance of almost all representations improved across all models, and the drop in the final layers of the discriminator is mitigated (cf. Figure 2(c) and Table 1 in Appendix D).
4.3 Parameter-Performance Trade-off
Finally, we investigate the parameter-performance trade-off. The results show the number of parameters the model has when using the best-performing representation on the validation set and the corresponding test set SCCs (cf. Figure 3 and Table 2). We consider only TMFT without improvements. Considering the test set scores and the number of parameters, the best models are BERT (tiny, mini, large) and ELECTRA generator (small, base, large). The difference in the number of parameters and layers suggests the depth of a transformer is essential for retaining performance on STS. Figure 3 also demonstrates an improvement in the number of parameters, and in the performance when comparing the last and best representation of the discriminator.
5 Conclusion
We analyze ELECTRA’s sentence embeddings for STS and two other tasks, comparing them to a BERT baseline on the STSB dataset. Our proposed truncated model fine-tuning method significantly improves the discriminator. The generator model matches BERT’s performance while significantly reducing the number of parameters and producing smaller embeddings.
6 Limitations
Our study reveals substantial differences in model performance on some tasks, but sentence embedding models can also be used for information retrieval. We limited ourselves to one dataset for STS, paraphrase identification, and entailment, four datasets for WS, and two datasets for MLM, due to computation restrictions. Experiments in Korean, German, and Spanish were exclusively conducted using the truncated model fine-tuning method. Another limitation of our work is the way we apply WS fine-tuning prior to truncated model fine-tuning on STS. We only consider fine-tuning the same layer’s embeddings for both procedures. Furthermore, in MLM, prior to truncated model fine-tuning on STS, we only use the last layer’s embeddings, as opposed to using previous layers. For all our experiments, we fixed the hyperparameters. Finally, our study covers only two families of models in detail: BERT and ELECTRA.
Appendix A Performance on Additional Tasks
Figure 4 represents the F1 test scores on the MRPC dataset. We use a bi-encoder optimized using a binary cross entropy loss.
Figure 5 represents the F1 test scores on the SICK entailment dataset. We use a bi-encoder to encode the first sentence , the second sentence , concatenate its absolute difference based on (Reimers and Gurevych, 2019). Finally, a classification head is used to calculate the output probabilities. The model is optimized using a cross entropy loss.
Appendix B Performance on STSB in Different Languages
The Korean STS (Ham et al., 2020) validation and test set use human error correction, while the other datasets are completely automatically translated using the DeepL API (May, 2021). The results for Korean, German, and Spanish are shown in Figures 6, 7, and 8 respectively. All experiments depict a performance drop in the final layers.
Appendix C Performance for Various Model Sizes on STSB
The figures 9, 10, and 11 show the test set SCCs on STSB across various model sizes. BERT and ELECTRA generator achieve the highest values in the last layer, while for the discriminator there is a visible performance drop.
Appendix D Truncated Model Fine-Tuning Results on English
D.1 Truncated Model Fine-Tuning and Improvements
Table 1 represents the test set SCCs and Pearson correlation coefficients on STSB for models that achieve the highest validation SCC using TMFT. The table also includes the results with the addition of improvements. Bold values represent the highest values for the used method.
Model Layer Valid Test TMFT BERT 12 86.07/85.98 82.74/83.03 ELECTRA 3 82.15/82.20 75.29/76.96 ELECTRA 12 86.62/86.38 82.57/82.50 DeBERTaV3 7 84.80/84.86 81.98/82.44 TMFT WS + TMFT STS BERT 12 85.85/85.84 83.77/84.08 ELECTRA 3 82.66/82.56 76.71/77.31 ELECTRA 12 86.67/86.39 82.85/82.91 MLM + TMFT STS BERT 12 85.91/85.71 82.66/82.64 ELECTRA 7 84.07/83.78 79.90/79.92 ELECTRA 11 85.58/85.22 80.97/80.85
D.2 Overview of Best-Performing Models Using TMFT on STSB
Table 2 represents the test set SCCs and Pearson correlation coefficients on STSB for models that achieve the highest validation SCC using TMFT. The table also includes the last included layer index and the number of parameters of the truncated model. Bold values represent the highest values for the model family.
Model | Layer | # Params (M) | Valid | Test |
BERT | 2 | 4.37 | 78.23/77.58 | 69.71/70.57 |
BERT | 4 | 11.10 | 83.10/82.47 | 75.64/76.36 |
BERT | 4 | 28.50 | 85.12/84.92 | 79.32/79.68 |
BERT | 8 | 41.11 | 85.72/85.39 | 80.84/81.10 |
BERT | 12 | 107.72 | 86.04/85.94 | 82.76/83.01 |
BERT | 24 | 332.53 | 88.27/88.25 | 85.58/85.75 |
ELECTRA | 1 | 4.76 | 79.72/79.26 | 68.82/69.63 |
ELECTRA | 12 | 13.45 | 74.02/73.15 | 66.54/67.00 |
ELECTRA | 3 | 45.10 | 82.14/82.18 | 75.29/76.96 |
ELECTRA | 12 | 108.89 | 72.32/71.56 | 66.93/67.35 |
ELECTRA | 12 | 182.94 | 84.36/84.40 | 80.54/80.71 |
ELECTRA | 24 | 334.09 | 35.96/32.66 | 35.53/33.09 |
ELECTRA | 12 | 13.45 | 84.67/84.15 | 81.79/81.16 |
ELECTRA | 12 | 33.31 | 86.60/86.36 | 82.54/82.48 |
ELECTRA | 24 | 50.74 | 87.14/86.72 | 84.61/84.47 |
Appendix E Representation Similarity on STSB
To explain the SCC decrease we provide a method using centered kernel alignment (CKA) (Kornblith et al., 2019). We apply the CKA to the STSB test set between representations obtained from the same hidden states of two models. Since the time complexity of the TMFT method we experiment with depends on the dataset size and the number of hidden states, it is beneficial to know about the presence of the drop prior to fine-tuning. An added bonus of this method is that it works with pre-trained models, saving time and resources. We use it as a proxy method for determining if the performance decrease will happen. We observe a pattern where the drop in CKA values in the final layers corresponds to the drop in SCC values in the final layers. The comparison between the ELECTRA discriminator and BERT (Figure 13) exhibits the same trend as the comparison between the DeBERTaV3 discriminator and BERT (Figure 14), which is a sharp decline in CKA values in the final few layers. The comparison between the ELECTRA generator and BERT (Figure 12) does not suffer from the CKA value drop. All combinations of inspected models after fine-tuning show similar trends in comparison to their pre-trained counterparts.
Appendix F Reproducibility
We conducted our experiments on an AMD Ryzen Threadripper 3970X 32-Core Processor and a single RTX 3090 GPU with 24GB of RAM. Running the experiments took around 270 GPU hours. DAPT experiments with BERT and ELECTRA discriminator take around 7 GPU hours, while for the ELECTRA generator it is around 4.5 GPU hours. The WS fine-tuning takes up to 3 minutes.